OpenStudy (anonymous):

Verify the identity. cot ( x - pi /2) = -tan x

OpenStudy (anonymous):

tan (pi / 2 - x) = cot x Which means cot (pi / 2 - x) = tan x since cot x = 1 / tan x Also tan (a - b) = -tan (b - a) And cot (a - b) = -cot (b - a)

OpenStudy (anonymous):

cot ( x - pi /2) = -tan x LHS = cot ( x - pi /2) = -cot ( pi /2-x)=-tanx =RHS Thus Proved

OpenStudy (anonymous):

What does this mean you all lost me explain it please

OpenStudy (anonymous):

tan (pi / 2 - x) = 1 / tan x = cot x 1 / tan (pi / 2 - x) = 1 / (1 / tan x) cot (pi / 2 - x) = tan x Next tan (a - b) = tan (-(b - a)) = - tan (b - a) 1 / tan (a - b) = -1 / tan (b - a) cot (a - b) = -cot (b - a)

OpenStudy (anonymous):

Where do you get tan ( a-b)

OpenStudy (anonymous):

Thats just a formula -.-

OpenStudy (anonymous):

Much better tan (-x) = -tan x

OpenStudy (anonymous):

Ok so could tell me how to work it out

OpenStudy (anonymous):

cot (x - pi / 2) = cot (-(pi / 2 - x)) = - cot (pi / 2 - x) = -tan x

OpenStudy (anonymous):

Ok so thats it pretty much